# An efficient wavelet-based physics-informed neural networks for singularly perturbed problems

### v1

@@ Abstract
Physics-informed neural networks (PINNs) are a class of deep learning models that utilize physics in the form of differential equations to address complex problems, including ones that may involve limited data availability. However, tackling solutions of differential equations with oscillations or singular perturbations and shock-like structures becomes challenging for PINNs. Considering these challenges, we designed an efficient wavelet-based PINNs (W-PINNs) model to solve singularly perturbed differential equations. Here, we represent the solution in wavelet space using a family of smooth-compactly supported wavelets. This framework represents the solution of a differential equation with significantly fewer degrees of freedom while still retaining in capturing, identifying, and analyzing the local structure of complex physical phenomena. The architecture allows the training process to search for a solution within wavelet space, making the process faster and more accurate. Further, the proposed model does not rely on automatic differentiations for derivatives involved in differential equations and does not require any prior information regarding the behavior of the solution, such as the location of abrupt features. Thus, through a strategic fusion of wavelets with PINNs, W-PINNs excel at capturing localized nonlinear information, making them well-suited for problems showing abrupt behavior in certain regions, such as singularly perturbed problems. The efficiency and accuracy of the proposed neural network model are demonstrated in various test problems, i.e., highly singularly perturbed nonlinear differential equations, the FitzHugh-Nagumo (FHN), and Predator-prey interaction models. The proposed design model exhibits impressive comparisons with traditional PINNs and the recently developed wavelet-based PINNs, which use wavelets as an activation function for solving nonlinear differential equations.
@@
Physics-informed neural networks, Wavelets, Singularly perturbed problems
@@ 1 Introduction
Singularly perturbed problems (SPPs) can be described using differential equations featuring at least one small parameter, popularly known as the “perturbation parameter(s)”. Specifically, when the parameter is very small (close to zero), the solution exhibits thin transition layers, often adjacent to the boundaries of the domain of interest. Consequently, the solution or its derivatives undergo rapid fluctuations within specific domain areas while maintaining smooth behavior outside these regions [1]. These problems are encountered in various physical systems that simultaneously contain both slow and fast varying components in the fields of science and engineering such as fluid dynamics [2, 3, 4], chemical reactors [5], financial modeling [6], gas porous electrodes [7], mathematical biology [8], modeling of semiconductor devices [9], and convective heat transfer problems with high Péclet numbers [10].
@@
The numerical solutions of these parameter-dependent problems based on the discretization approaches are impractical since it is affected by the parameter. Further, a large number of mesh points are required to produce sufficiently precise approximations of the actual solution for various perturbation parameter values. Many methods have been developed to address this challenge, such as parameter-uniform, parameter-robust, or uniformly convergent approaches like Fitted Operator Methods (FOMs) [11, 12, 13] and Fitted Mesh Methods (FMMs) [14, 15]. Moreover, to address the challenges with meshes, various meshfree methods, such as the moving least squares (MLS) method [16], element-free Galerkin approach [17], and the local Petrov–Galerkin method [18], have been developed to solve SPPs numerically [19, 20].
@@
This paper concentrates on one such class of methods, which eliminates not only meshes but also has the ability to handle complex geometry and higher-dimensional problems, namely, physics-informed neural networks (PINNs). A few of the earliest works on solving differential equations utilizing neural networks were presented by Dissanayake et al. [21] and Lagaris et [22]. Their method required a trial solution that satisfies all boundaries and initial conditions. However, this technique often fails if a problem has complex behavior, and finding appropriate trial solutions for such problems becomes challenging. Later, Raissi et al. [23] introduced PINNs, which incorporate initial and boundary conditions into the network’s loss function without considering a trial solution beforehand. Following this work, different types of PINNs, such as variational physics-informed neural networks (VPINN) [24], gradient-enhanced physics-informed neural networks (GPINN) [25], and extended physics-informed neural networks (XPINN) [26], enhanced physics-informed neural networks with Augmented Lagrangian relaxation method (AL-PINNs) [27], and physics-informed extreme learning machines [28, 29], have been developed to enhance the performance of conventional PINN. For more details, see a comprehensive review of PINNs [30, 31] and references therein.
@@
Despite many significant advantages, PINNs encounter challenges in accurately handling equations with rapid oscillations [31]. This difficulty arises because the various components in the loss function have varying convergence rates during training. PINNs need to efficiently represent and learn these features that vary significantly across various scales. While existing theories address issues related to isolated rapid changes, using techniques like domain splitting [32], these approaches can be cumbersome to implement. Furthermore, the authors of [33] introduced an improved physics-informed neural networks algorithm, which involved some modifications from the approach described in [32]. A singularity term was included in the calculation formula to account for the singularity of the equation. Other approaches [34, 35] in which authors sampled dense training points near high-gradient regions to improve convergence. Motivated by the reviewed research and identified shortcomings, this work proposes a novel modified version of PINNs i.e., wavelet-based physics-informed neural networks (W-PINNs), which exploit the ability of wavelets to address these challenges.
@@
The localization properties of wavelets make them beneficial when analyzing problems with localized features, such as initial or boundary layers, in SPPs. Mallat’s work [36] and Daubechies’ contributions [37] lay down the foundational principles of wavelet analysis. Using the Haar wavelet method, Kumar and Deswal [38] have developed two-parameter SPPs with Robin boundary conditions. Further, Kumar et al. designed the Haar wavelet-based approximation scheme for singularly perturbed partial integro-differential equations [39]. Then, Ahsan et al. introduced a wavelet collocation method for solving SPPs with integral boundary conditions [40]. Wavelets excel at capturing nonlinear information within localized regions of the solution. In contrast, neural networks are well-known for learning and modeling global nonlinear relationships. By incorporating wavelets with neural networks, we can potentially leverage the ability of wavelets to extract these localized nonlinearities. This can complement the neural network’s nonlinear learning capability and potentially enhance the overall ability to capture intricate nonlinear patterns, possibly improving the accuracy of the network. The incorporation of wavelets into PINNs has been attempted in [41]. However, this paper focuses on wavelets solely as activation functions, with the overall structure of PINNs remains unchanged from its original introduction. Our findings show that this method also fails to approximate SPPs satisfactorily.
@@
In this study, we use the wavelet family to represent an approximate solution to the differential equation, and the weight of each family member is optimized using the W-PINN. Unlike previous neural network-based approaches for solving SPPs, this method requires no prior information about the nature of the solution, making it practical and easy to implement. Additionally, the proposed neural architecture significantly reduces training time as it does not rely on automatic differentiation for the derivatives involved in the loss function. We use Mexican hat and Gaussian wavelets for implementation, and their performance is compared with conventional PINNs to validate the efficacy. Our approach demonstrates high accuracy across various singularly perturbed ordinary differential equations examined in this research and establishes it as a robust method for these class problems.
@@
The organization of the remaining paper is outlined as follows: Section 2 initiates with the introduction of standard PINNs, followed by a thorough discussion on the design and operational mechanism of W-PINNs. Section 3 presents numerical results for various SPPs to demonstrate the effectiveness and accuracy of the proposed method. Finally, Section 4 serves as the concluding segment, offering a concise summary of the key findings, contributions, and future work along with limitations.
@@ 2.1 Physics-informed neural networks (PINNs)
where $\boldsymbol{z}^{k}\in\mathds{R}^{n_{k}}$ is the outcome of $k$-th layer, $\boldsymbol{\omega}^{k}\in\mathds{R}^{n_{k}\times n_{k-1}}$, $\boldsymbol{b}^{k}\in\mathds{R}^{n_{k}}$, and $n_{k}$ is the number of neurons in $k^{th}$ layer. Here, $\boldsymbol{\omega}$ and $\boldsymbol{b}$ denote the weight and bias, the model parameters to be optimized. Moreover, $\sigma$ is a nonlinear mapping known as an activation function. This process is widely known as feed-forward propagation. We define a loss function ($\mathcal{L}$) for network output, which quantifies the disparity between the network’s predictions and the desired outcomes. The neural network is trained using backpropagation [43], a technique wherein optimization algorithms, such as gradient descent, to minimize loss function and fine-tune the model parameters ($\boldsymbol{\omega},\boldsymbol{b}$). Trained parameters capture the underlying structure of the problem in consideration.
Figure 1: A PINN architecture for ODEs.
It is well known that PINNs incorporate the underlying physics of the system by using a governing set of equations along with initial and boundary conditions into the loss function. It measures how much the network’s predicted solution violates the differential equation at several collocation points. Mathematically, it can be written as $\mathcal{L}=\mathcal{L}_{r}+\mathcal{L}_{c}$, where $\mathcal{L}_{r}$ denotes the residual’s mean squared error and $\mathcal{L}_{c}$ denotes the initial and boundary condition’s mean squared error. During optimization, a derivative of the loss function with respect to the model parameters and derivatives of the network’s output with respect to input parameters, exhibited in the loss function are required. These derivatives are effectively computed using “automatic differentiation” [44]. Thus, the objective of PINNs is to identify the optimal parameters that minimize the specified loss function, which encapsulates all physics-related information. Consider a general ordinary differential equation (ODE):
$\begin{cases}\mathscr{L}\big{(}x,u,u^{(1)},u^{(2)},\cdots,u^{(n)}\big{)}=0,~{}~{}x\in(a,b),&\\ \mathscr{B}\big{(}x(a),u(a),u^{(1)}(a),\cdots u^{(n-1)}(a),x(b),u(b),u^{(1)}(b),\cdots u^{(n-1)}(b)\big{)}=0,&\end{cases}$ (2)
where $\mathscr{L}$ denotes a differential operator and $\mathscr{B}$ represents initial and/or boundary conditions. To solve such ODEs, a fully connected schematic PINN architecture is shown in Fig (1).
@@ 2.2 Wavelet-based physics-informed neural networks (W-PINNs)
We start with a family of wavelets, whose members are dilates and translates of the mother wavelet. Indeed, the dilated and translated version of the mother wavelet function, $\psi(x)$ for $j,k\in\mathbb{Z}$ can be written as
@@
Here $j$ is the exponent of dilation, and $k$ is the translation parameter. In order to define the number of family members, we fix the maximum resolution for the family as $J$. For a given $J$, the dilation exponent, $j$, ranges from 00 to $J-1$, and for a $j,k$ ranges from 00 to $12^{j}-$ 1. This way of constructing a family enables us to cover the whole domain, $[0,1]$ with various resolutions. For instance, $J=2$ will consist of three members: $\psi_{0,0}(x),\psi_{1,0}(x),\psi_{1,1}(x)$.
Figure 2: A W-PINN architecture for second order ODEs.
Figure 3: A W-PINN architecture for coupled second order ODEs.
After constructing a family of wavelets, we approximate the solution of the differential equation as:
$\hat{u}(x)=\sum_{j=0}^{J-1}\sum_{k=0}^{2^{j}-1}c_{j,k}\psi_{j,k}(x)+\mathcal{B}~{}~{}\text{equivalently,}~{}~{}\hat{u}(x)=\sum_{i=0}^{2^{J}-1}c_{i}\psi_{i}+\mathcal{B},$ (3)
where $\mathcal{B}$ is bias, $i=2^{j}+k-1$ and $c_{i}$ denotes the coefficient of the $i^{th}$ member. In this study, we consider two mother wavelets, namely, the Gaussian and the Mexican Hat, which are single and double derivatives of the Gaussian function, respectively. The mathematical expressions for the Mexican hat $(\psi^{M}(x))$, Gaussian $(\psi^{G}(x))$ can be written as
$\psi^{M}(x)=(1-x^{2})e^{-\frac{x^{2}}{2}},~{}~{}~{}~{}\psi^{G}(x)=-xe^{-\frac{x^{2}}{2}}$
@@
Here, we added a bias $\mathcal{B}$ to make a degree of freedom in approximating solutions with non-zero means, as the Mexican Hat and Gaussian are symmetrical wavelets. This enables the model to search for a solution in a space that is more suitable for singularly perturbed problems.
We employ W-PINN to obtain an approximate solution, as described above. W-PINN is composed of two parts: a trainable network that predicts coefficients by forming a fully connected network, and a non-trainable network that provides the final approximated solution and its derivatives. Once we have an approximate solution, we get the loss function of the network using the mean square of residuals and the mean square error of the initial and boundary conditions. The schematic diagrams of W-PINN for second-order ODEs and coupled ODEs are shown in Figure 2 and Figure 3, respectively. W-PINN requires collocation points spanning all over the domain as a training set and a wavelet matrix and its derivative matrices up to the order of the differential equation in consideration. These wavelet matrices serve as the weights of the non-trainable network. A trainable bias is also added to the wavelet matrix $(\mathbf{W})$ in order to incorporate $\mathcal{B}$. For a given set of collocation points $\mathbf{x}=(x_{0},x_{1},\cdots,x_{N})$, the wavelet matrices are constructed as follows:

### v2

@@ Abstract
Physics-informed neural networks (PINNs) are a class of deep learning models that utilize physics in the form of differential equations to address complex problems, including ones that may involve limited data availability. However, tackling solutions of differential equations with rapid oscillations, steep gradients, or singular behavior becomes challenging for PINNs. Considering these challenges, we designed an efficient wavelet-based PINNs (W-PINNs) model to address this class of differential equations. Here, we represent the solution in wavelet space using a family of smooth-compactly supported wavelets. This framework represents the solution of a differential equation with significantly fewer degrees of freedom while still retaining the dynamics of complex physical phenomena. The architecture allows the training process to search for a solution within the wavelet space, making the process faster and more accurate. Further, the proposed model does not rely on automatic differentiations for derivatives involved in differential equations and does not require any prior information regarding the behavior of the solution, such as the location of abrupt features. Thus, through a strategic fusion of wavelets with PINNs, W-PINNs excel at capturing localized nonlinear information, making them well-suited for problems showing abrupt behavior in certain regions, such as singularly perturbed and multiscale problems. The efficiency and accuracy of the proposed neural network model are demonstrated in various 1D and 2D test problems, i.e., the FitzHugh-Nagumo (FHN) model, the Helmholtz equation, the Maxwell equation, and the Allen-Cahn equation, along with other highly singularly perturbed nonlinear differential equations. The proposed model significantly improves with traditional PINNs, recently developed wavelet-based PINNs, and other state-of-the-art methods.
@@
Physics-informed neural networks, Wavelets, Singularly perturbed problems, Multiscale problems
@@ 1 Introduction
Singularly perturbed problems (SPP.s) can be described using differential equations featuring at least one small parameter, popularly known as the “perturbation parameter(s)”. Specifically, when the parameter is very small (close to zero), the solution exhibits thin transition layers, often adjacent to the boundaries of the domain of interest. Consequently, the solution or its derivatives undergo rapid fluctuations within specific domain areas while maintaining smooth behavior outside these regions [1]. These problems are encountered in various physical systems that simultaneously contain both slow and fast varying components in the fields of science and engineering such as fluid dynamics [2], chemical reactors [3], financial modeling [4], modeling of semiconductor devices [5], and convective heat transfer problems with high Péclet numbers [6].
@@
The numerical solutions of these parameter-dependent problems based on the discretization approaches are impractical since it is affected by the parameter. Further, a large number of mesh points are required to produce sufficiently precise approximations of the actual solution for various perturbation parameter values. Many methods have been developed to address this challenge, such as parameter-uniform, parameter-robust, or uniformly convergent approaches like Fitted Operator Methods (FOMs) [7, 8] and Fitted Mesh Methods (FMMs) [9, 10]. Moreover, to address the challenges with meshes, various meshfree methods, such as the moving least squares (MLS) method [11], element-free Galerkin approach [12], and the local Petrov–Galerkin method [13], have been developed to solve SPPs numerically.
@@
This study concentrates on one such class of methods, which not only eliminates meshes but also has the ability to handle complex geometry and higher-dimensional problems, namely, physics-informed neural networks (PINNs). A few of the earliest works on solving differential equations utilizing neural networks were presented by Dissanayake et al. [14] and Lagaris et [15]. Their method required a trial solution that satisfies all boundaries and initial conditions. However, this technique often fails if a problem has complex behavior, and finding appropriate trial solutions for such problems becomes challenging. Later, Raissi et al. [16] introduced PINNs, which incorporate initial and boundary conditions into the network’s loss function without considering a trial solution beforehand. Following this work, different types of PINNs, such as variational physics-informed neural networks (VPINN) [17], gradient-enhanced physics-informed neural networks (GPINN) [18], and extended physics-informed neural networks (XPINN) [19] and many other have been developed to enhance the performance of conventional PINN. For more details, see a comprehensive review of PINNs [20, 21] and references therein.
@@
Despite many significant advantages, PINNs encounter challenges in accurately handling equations with sharp gradients or rapid oscillations [21]. This difficulty arises because the various components in the loss function have varying convergence rates during training. PINNs need to efficiently represent and learn these features that vary significantly across various scales. While existing theories address issues related to isolated rapid changes, using techniques like domain splitting [22, 23] or using trainable weights to balance loss terms dynamically [24], these approaches can be cumbersome to implement. Furthermore, the authors of [25] introduced an improved physics-informed neural networks algorithm, which involved some modifications from the approach described in [22]. A singularity term was included in the calculation formula to account for the singularity of the equation. Other approaches [26, 27] in which authors sampled dense training points near high-gradient regions to improve convergence. Motivated by the reviewed research and identified shortcomings, this work proposes a novel modified version of PINNs i.e., wavelet-based physics-informed neural networks (W-PINNs), which exploit the ability of wavelets to address these challenges.
@@
The localization properties of wavelets make them beneficial when analyzing problems with localized features, such as initial or boundary layers, in SPPs [28, 29]. Mallat’s work [30] and Daubechies’ contributions [31] lay down the foundational principles of wavelet analysis. Wavelets excel at capturing nonlinear information within localized regions of the solution [32, 33]. In contrast, neural networks are well-known for learning and modeling global nonlinear relationships. By incorporating wavelets with neural networks, we can potentially leverage the ability of wavelets to extract these localized nonlinearities. This can complement the neural network’s nonlinear learning capability and potentially enhance the overall ability to capture intricate nonlinear patterns, possibly improving the accuracy of the network. The incorporation of wavelets into PINNs has been attempted in [34]. However, the authors of [34] focused on wavelets solely as activation functions, with the overall structure of PINNs remains unchanged from its original introduction. Our findings show that this method also fails to approximate SPPs satisfactorily.
@@
In this study, we use the wavelet family to represent an approximate solution to the differential equation, and the weight of each family member is optimized using the W-PINN. Unlike previous neural network-based approaches, the proposed neural architecture does not rely on automatic differentiation for the derivatives involved in the loss function, which significantly reduces training time. Additionally, this method requires no prior information about the nature of the solution, making it practical and easy to implement. We use Gaussian and Mexican hat wavelets for implementation, and their performance is compared with conventional PINN and state-of-the-art methods to validate the efficacy. Our approach demonstrates high accuracy across a range of differential equations exhibiting steep gradients, rapid oscillations, singularities, and multiscale behavior, establishing it as a robust method for these classes of problems. The key contributions of this work can be summarized as follows:
  • W-PINN introduces a framework that eliminates automatic differentiation in loss function derivative computation, significantly accelerating training while maintaining or improving solution accuracy compared to recent methods in the literature.
  • W-PINN effectively addresses the loss balancing challenges inherent in conventional PINNs, particularly for problems exhibiting multi-scale phenomena, singular behavior, or rapid oscillations in their solutions.
  • The proposed method is validated through the FHN model, Helmholtz equation, Allen-Cahn equation, Maxwell equation, Lid-driven cavity flow, and various other singularly perturbed problems.
@@
The organization of the remaining paper is outlined as follows: Section 2 initiates with the introduction of standard PINNs, followed by a thorough discussion on the design and operational mechanism of W-PINNs. Section 3 presents numerical results for various differential equations to demonstrate the effectiveness and accuracy of the proposed method. Finally, Section 4 serves as the concluding segment, offering a concise summary of the key findings, contributions, and future work along with limitations.
@@ 2.1 Physics-informed neural networks (PINNs)
where $\boldsymbol{z}^{k}\in\mathds{R}^{n_{k}}$ is the outcome of $k$-th layer, $\boldsymbol{\omega}^{k}\in\mathds{R}^{n_{k}\times n_{k-1}}$, $\boldsymbol{b}^{k}\in\mathds{R}^{n_{k}}$, and $n_{k}$ is the number of neurons in $k^{th}$ layer. Here, $\boldsymbol{\omega}$ and $\boldsymbol{b}$ denote the weight and bias, the model parameters to be optimized. Moreover, $\sigma$ is a nonlinear mapping known as an activation function. This process is widely known as feed-forward propagation. We define a loss function ($\mathcal{L}$) for network output, which quantifies the disparity between the network’s predictions and the desired outcomes. The neural network is trained using backpropagation [36], a technique wherein optimization algorithms, such as gradient descent, to minimize loss function and fine-tune the model parameters ($\boldsymbol{\omega},\boldsymbol{b}$). Trained parameters capture the underlying structure of the problem in consideration. A well-known class of neural network is physics-informed neural networks (PINNs), which incorporate the underlying physics of the system by using a governing set of equations along with initial and boundary conditions into the loss function. It measures how much the network’s predicted solution violates the differential equation at several collocation points. Mathematically, it can be written as $\mathcal{L}=\mathcal{L}_{\text{res}}+\mathcal{L}_{\text{ic}}+\mathcal{L}_{\text{bc}}$, where $\mathcal{L}_{\text{res}}$ denotes the residual mean squared error, $\mathcal{L}_{\text{ic}}$ and $\mathcal{L}_{\text{bc}}$ denote mean squared errors for initial and boundary conditions respectively. During optimization, the derivative of the loss function with respect to the model parameters and derivatives of the network’s output with respect to input parameters, exhibited in the loss function are required. These derivatives are effectively computed using “automatic differentiation” [37]. Thus, the objective of PINNs is to identify the optimal parameters that minimize the specified loss function, which encapsulates all physics-related information.
@@ 2.2 Wavelet-based physics-informed neural networks (W-PINNs)
PINNs have emerged as a promising mesh-free approach for solving wide varieties of differential equations. However, their performance significantly deteriorates when dealing with problems exhibiting high gradients, rapid oscillations or singular behavior, primarily due to challenges in balancing different loss terms and accurately computing derivatives through automatic differentiation. Recent literature has proposed several approaches to address these limitations. McClenny et. al. introduced Self-adaptive PINNs (SA-PINNs)[24] that employ trainable weights to balance loss terms dynamically, while Wang et al. (2024) developed a practical PINN framework for multi-scale problems with multi-magnitude loss terms (MMPINN)[23] framework incorporating specialized architectures and regularization strategies for multi-scale problems. While these methods demonstrate improved performance, they still face computational overhead from automatic differentiation and require careful tuning of multiple hyperparameters. The proposed wavelet-based PINN (W-PINN) framework offers an elegant solution to these challenges by expressing the solution as a linear combination of wavelet basis functions. By training the coefficients rather than the solution directly, W-PINN eliminates the need for automatic differentiation in computing derivatives in loss function, as they can be obtained analytically from the wavelet basis functions. This approach not only significantly reduces computational cost but also naturally handles multi-scale features and balances different loss terms.
Formulation of W-PINN starts with a family of wavelets, whose members are dilates and translates of the mother wavelet. The dilated and translated version of the mother wavelet function, $\psi(x)$ for $j,k\in\mathbb{Z}$ can be written as
@@
[…]
$\hat{u}(x)=\sum_{j=J_{1}}^{J_{N}}~{}\sum_{k=\lceil a\cdot 2^{j+1}\rceil}^{\lceil b\cdot 2^{j+1}\rceil}c_{j,k}\Psi_{j,k}(x)+\mathcal{B},$ (2)
where $\mathcal{B}$ is bias, $J_{i}$ is $i^{th}$ resolution and $N$ is the total number of resolutions in set $J$. Extension of this formulation for higher dimensional problem is straight forward, for instance here is formulation for a 2D problem:
$\Psi_{j_{1},j_{2},k_{1},k_{2}}(x,y)$ $=\sqrt{2^{j_{1}}\cdot 2^{j_{2}}}\psi_{X}(2^{j_{1}}x-k_{1})~{}\psi_{Y}(2^{j_{2}}y-k_{2}),$ (3)
$\hat{u}(x,y)$ $=\sum_{j_{1}=J_{11}}^{J_{1N_{1}}}~{}\sum_{j_{2}=J_{21}}^{J_{2N_{2}}}~{}\sum_{k_{1}=\lceil a_{1}\cdot 2^{j_{1}+1}\rceil}^{\lceil b_{1}\cdot 2^{j_{1}+1}\rceil}~{}\sum_{k_{2}=\lceil a_{2}\cdot 2^{j_{2}+1}\rceil}^{\lceil b_{2}\cdot 2^{j_{2}+1}\rceil}c_{j_{1},j_{2},k_{1},k_{2}}\Psi_{j_{1},j_{2},k_{1},k_{2}}(x,y)+\mathcal{B},$ (4)
where the domain $\Omega=[a_{1},b_{1}]\times[a_{2},b_{2}]$ is discretized using resolution parameters $J_{1}$ and $J_{2}$ corresponding to the $x$ and $y$ dimensions respectively. The translation parameters $k_{1}$ and $k_{2}$ define the positioning of basis functions along these respective dimensions. This hierarchical construction of wavelet families at different resolutions enables capturing solution features at multiple scales - from broad, global behavior to fine, local details. In this study, we consider two mother wavelets, namely, the Gaussian and the Mexican Hat, which are single and double derivatives of the Gaussian function, respectively. The mathematical expressions for the Mexican hat $(\psi^{M}(x))$, Gaussian $(\psi^{G}(x))$ can be written as
$\psi^{M}(x)=(1-x^{2})e^{-\frac{x^{2}}{2}},~{}~{}~{}~{}\psi^{G}(x)=-xe^{-\frac{x^{2}}{2}}.$
@@
In equations (2) and (4), we added a bias $\mathcal{B}$ to make a degree of freedom in approximating solutions with non-zero means, as the Mexican Hat and Gaussian are symmetrical wavelets.
Figure 1: A W-PINN architecture.
We employ W-PINN to obtain an approximate solution, as described above. W-PINN is composed of two parts: a trainable network that predicts coefficients by forming a fully connected network, and a non-trainable network that constructs the final solution and its derivatives using these coefficients and pre-computed wavelet matrices. This formulation eliminates the need for automatic differentiation (AD) in computing PDE residuals. While AD provides machine-precision derivatives, but it comes at cost of maintaining computation-intensive computational graphs that grow increasingly complex with network depth and order of derivatives involved. Our wavelet-based approach not only makes training faster by eliminating need of computational graph but also avoid potential numerical instabilities that commonly arise in AD during training, particularly in problems exhibiting high gradients or singular behavior. This approach aligns with recent developments in the field, such as DT-PINN [38], which achieved two to four times speedups over AD-based PINNs by eliminating AD through meshless radial basis function-finite differences (RBF-FD). However, our wavelet-based formulation provides the additional advantage of naturally handling multi-scale features through its hierarchical basis structure. The architecture of W-PINN for a two-dimensional problem is illustrated in Figure 1. In general, for a $d$-dimensional problem, the network architecture comprises an input layer with $d$ neurons corresponding to the spatial-temporal dimensions. These inputs are processed through a shallow network that maps each collocation point to a corresponding entry in the feature layer. For one-dimensional problems, this mapping can be simplified by directly using the input layer as the feature layer. Now, the feature layer passes through a fully connected network to get coefficients. To further enhance model performance, these coefficients can be separately fine-tuned through hyperparameter optimization. The final solution is then computed by combining these optimized coefficients with the pre-calculated wavelet matrices. Once we have an approximate solution and its derivatives, we get the loss function of the network using the mean square of residuals and the mean square error of the initial and boundary conditions. W-PINN requires collocation points spanning all over the domain as a training set and a wavelet matrix and its derivative matrices up to the order of the differential equation in consideration. These wavelet matrices serve as the weights of the non-trainable network. A trainable bias is also added to the wavelet matrix $(\mathbf{W})$ in order to incorporate $\mathcal{B}$. For a given set of collocation points $\mathbf{x}=(x_{0},x_{1},\cdots,x_{N})$, the wavelet matrices are constructed as follows: